# font=small

### v1

@@ 2 Short-Cavity Regime
An alternative and trivial means to reproduce the Rabi oscillation frequency is through a rescaling of the atomic TDM, $\mu_{12}$, while fixing the mode occupancy to its physically-informed value, $n^{(0)}=1/2$. For DC-MF dynamics, the applicable rescaling factor amounts to 1.33, whereas a value of 1.06 applies to MF dynamics. Conceptually, such a rescaling points to an effective underestimation of the dipolar effect under the applied mixed quantum–classical schemes. In the following, we will invoke this rescaling approach while applying focused sampling to MF and DC-MF dynamics, which henceforth will be referred to as MF(f1.06) and DC-MF(f1.33) dynamics, respectively. Our particular choice of rescaling factors is affirmed by Fig. 2 (b), which presents results from MF(f1.06) and DC-MF(f1.33) dynamics showing favorable agreement with CISD results, although we note the minor beating pattern exhibited by DC-MF(f1.33) dynamics due to the presence of low-frequency anharmonicity. Evidently, focused sampling is seen to effectively resolve the spurious destructive interference caused by Wigner sampling. (It should be noted that the value of $g$ is kept unchanged in this analysis, meaning that $g$ is continually expressed in terms of the unscaled TMDs.)
@@ 3 Long-Cavity Regime
Having established that DC-MF(f1.33) dynamics yields desirable results in the short-cavity limit, we proceed to revisit the long-cavity regime for which DC-MF dynamics was originally developed 36, 37. Under Wigner sampling, DC-MF dynamics was previously shown to reach near-quantitative accuracy in this regime 36, 37, while MF dynamics proved broadly capable of reproducing qualitative trends 34, 35. It would thus be interesting to make a comparative assessment of the performance of both methods under focused sampling.
As in our previous evaluations of the long-cavity regime 36, 37, we parametrize our atom–cavity system following the example of Hoffmann et al. 34, such that the atom is taken to represent the lowest two levels of a one-dimensional hydrogen model 50, 43, with energies $\epsilon_{1}=-0.6738$ a.u. and $\epsilon_{2}=-0.2798$ a.u., and with a TDM $\mu_{12}=1.034$ a.u., whereas the cavity length is taken as $L=2.362\times 10^{5}~{}\mathrm{a.u.}~{}=12.5~{}\mathrm{\mu m}$. The 400 lowest-energy cavity modes are incorporated in our simulations, which suffices to ensure convergence.
@@
Before presenting any graphical results, we report that a simple switching from Wigner sampling to focused sampling does not appreciably change the results generated by MF and DC-MF dynamics for this long-cavity implementation. We consider this a desirable outcome, since DC-MF(W) dynamics did already perform desirably to begin with 36, although one could have wished that the change in sampling approach would resolve inaccuracies seen for MF(W) dynamics 34, 36, 37. Regardless, such switching from Wigner sampling to focused sampling allows the number of trajectories, necessary for convergence, to be reduced from (roughly) 500 to 30 for DC-MF dynamics, while a reduction from 5000 to 2500 was found for MF dynamics.
With regard to modeling accuracy, things turn interesting once we apply the aforementioned TDM rescaling to both methodologies. Fig. 4 compares DC-MF(f1.33) dynamics to DC-MF(W) dynamics where, remarkably, the TDM rescaling is seen to almost fully remedy the lag known 36, 37 to occur for the atomic excited-state occupancy within DC-MF(W) dynamics. As such, the TDM rescaling yields a further-improved agreement of this method with CISD results, with minor discrepancies occurring exclusively in the reabsorption events. Also included in the comparison is MF(f1.06) dynamics, results of which show qualitative behavior consistent with previous MF(W) dynamics simulations 34, 35.
Figure 5: Optical field intensity at $t=500$ a.u. (a), $t=2200$ a.u. (b), and $t=4000$ a.u. (c), which correspond to instances shortly after the first emission, the first reabsorption, and the second reabsorption events, respectively. Note that the optical field is symmetric with respect to reflection through the cavity center, and only the right half of the field intensity is depicted. Shown are results from DC-MF(W) (blue), MF(f1.06) (orange), and DC-MF(f1.33) (purple) dynamics, compared against those from CISD (gray). Each trace was smoothed by a $0.125$ $\mathrm{\mu}$m moving average for ease of demonstration.
@@
As with the atomic occupancies, optical field intensity profiles generated by MF dynamics and DC-MF dynamics are invariant with regard to whether Wigner sampling or focused sampling is applied, but TDM rescaling has a significant effect on DC-MF dynamics. Shown in Fig. 5 are optical field intensity results for DC-MF(W), DC-MF(f1.33), and MF(f1.06) dynamics, compared against CISD results at different time instances. While DC-MF(W) dynamics captures most qualitative features seen in the CISD results, as discussed before 34, 36, it is again remarkable to see the extent to which TDM rescaling improves the accuracy of this method to a near-quantitative level. Indeed, DC-MF(f1,33) dynamics reproduces subtle details of the lagging wavefronts that arise upon reabsorption 36, with differences creeping in only after two reabsorption events. Combined with the favorable results obtained for the half-wavelength cavity, DC-MF(f1.33) dynamics thus presents itself as a method with applicability across cavity lengths that relies on invariant TDM rescaling factors.
@@ 4 Conclusions and Outlook
In conclusion, DC-MF dynamics offers a route to accurate and low-cost simulations of cQED across cavity lengths, provided that focused sampling and TDM rescaling are invoked. In demonstrating this, we built heavily on recent work 38 where MF dynamics (without the decoupling scheme) was shown to reproduce full-quantum results for a half-wavelength cavity. Such was achieved under a focused sampling of initial optical field coordinates, in combination with a slight renormalization of the initial optical ZPE. Here, we present TDM rescaling as an alternative renormalization procedure that not only produces equally accurate results for MF dynamics, but that also allows DC-MF dynamics to produce desirable results. As mentioned, we find that switching from the commonly-used Wigner sampling technique to focused sampling has only a minor effect on Ehrenfest-based mixed quantum–classical modeling in the long-cavity regime, meaning that it fortuitously does not compromise the high accuracy that DC-MF dynamics is known to reach 36, 37. TDM rescaling, on the other hand, provides a further source of accuracy improvement for this method, yielding atomic occupancies and optical field intensity profiles in excellent agreement with full-quantum results.
@@
We consistently find focused sampling to dramatically improve the accuracy of MF and DC-MF dynamics in the short-cavity regime, yet previous applications of this sampling technique were primarily concerned with a reduction of computational cost 40, 41, 48, 47, 44, 45, 46. While we did not systematically study the computational cost savings afforded by focused sampling in cQED, which we reserve for a future study, we anecdotally found such savings to oftentimes be significant, especially in the short-cavity regime, where a reduction of the necessary number of trajectories was found of three orders of magnitude. In the long-cavity regime, this reduction was much more modest, with a factor of two being typical. Moreover, we find the number of trajectories necessary for convergence to be strongly dependent on the observable under consideration. Notably, transient occupancies of quantum emitters require about 10 to 30 trajectories for DC-MF dynamics under focused sampling, a number much lower than that typically used in mixed quantum–classical simulations of population dynamics. For the optical field intensity, this number raises to values in the range 50 000 to 100 000, which are more typical.
@@
The results presented in this Article help to further establish DC-MF dynamics as a mixed quantum–classical approach for cQED simulations that optimally balances accuracy and computational efficiency. This not only promotes the use of DC-MF dynamics in studying cQED based on model systems, but it also bears relevance to a broader effort to incorporate quantum modeling in established techniques from classical optics 51, 52, 53, 54, 55, 56, including finite-difference time-domain, for which DC-MF dynamics offers a preferred route. The cost associated with trajectory sampling is very modest, so long as transient matter-based occupancies are being considered, while the evaluation of optical field intensity profiles remains computationally much more challenging.
@@
Against the favorable performance of DC-MF dynamics, there are various aspects of this method that await a deeper rationale. One such aspect is concerned with the form of the decoupling term appearing in Eq. 10. Broadly speaking, this term may be considered representative of a functional of the atomic quantum state, which vanishes when the ground state is occupied while reaching unity for a fully-occupied excited state 36. Future efforts may unveil the existence of an exact functional, or at least identify improved proxies that further enhance the method’s accuracy. Another aspect that deserves further attention is concerned with the TDM rescaling factors introduced in the present Article. Ultimately, these factors compensate for the tendency of mixed quantum–classical modeling to overshoot the optical occupancy needed to reproduce Rabi oscillations, which in turn is due to its oscillation frequency decaying to zero in the limit $n^{(0)}\rightarrow 0$ 38. It would be interesting to further investigate this behavior, possibly through an analysis similar to that presented in Ref. 38. Altogether, such inquiries may provide further guidance for the application of mixed quantum–classical dynamics to the modeling of transient phenomena.

### v3

@@ 3 Short-Cavity Regime
In the following, we will invoke this rescaling approach while applying focused sampling to MF and DC-MF dynamics, which henceforth will be referred to as MF(f1.06) and DC-MF(f1.33) dynamics, respectively. Our particular choice of rescaling factors is affirmed by Fig. 4, which presents results from MF(f1.06) and DC-MF(f1.33) dynamics showing favorable agreement with CISD results, although we note the minor beating pattern exhibited by DC-MF(f1.33) dynamics due to the presence of low-frequency anharmonicity. Evidently, focused sampling is seen to effectively resolve the spurious destructive interference caused by Wigner sampling for both methods, and the renormalization of TDMs in the light–matter interaction Hamiltonian helps reproduce the correct Rabi frequency. (Analogous results for MF and DC-MF dynamics under focused sampling but without TDM renormalizations are presented in Fig. S1 of the SI, showing persistent Rabi oscillations, but at an underestimated frequency.)
@@ 4 Long-Cavity Regime
Having established that MF(1.06) and DC-MF(f1.33) dynamics yield desirable results in the short-cavity limit, we proceed to revisit the long-cavity regime for which DC-MF dynamics was originally developed hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. Under Wigner sampling, DC-MF dynamics was previously shown to reach near-quantitative accuracy in this regime hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023, while MF dynamics proved broadly capable of reproducing qualitative trends hoffmannCapturingVacuumFluctuations2019, hoffmannBenchmarkingSemiclassicalPerturbative2019. It would thus be interesting to make a comparative assessment of the performance of both methods under focused sampling and TDM renormalizations.
As in our previous evaluations of the long-cavity regime hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023, we parametrize our atom–cavity system following the example of Hoffmann et al. hoffmannCapturingVacuumFluctuations2019, such that the atom is taken to represent the lowest two levels of a one-dimensional hydrogen model suModelAtomMultiphoton1991, flickAtomsMoleculesCavities2017, with energies $\epsilon_{1}=-0.6738$ a.u. and $\epsilon_{2}=-0.2798$ a.u., and with a TDM $\mu_{12}=1.034$ a.u., whereas the cavity length is taken as $L=2.362\times 10^{5}~\mathrm{a.u.}~=12.5~\mathrm{\mu m}$. The 400 lowest-energy cavity modes are incorporated in our simulations, which suffices to ensure convergence.
@@
First, we report that focusing the initial cavity mode occupancies as $n^{(0)}=1/2$ yields results virtually indistinguishable from those from Wigner sampling for both MF and DC-MF dynamics (see Fig. S2 in the SI). This implies that the destructive interference seen for Rabi oscillations in the half-wavelength cavity does not negatively affect the long-cavity limit. We consider this a desirable outcome, since DC-MF(W) dynamics did already perform favorably for this long-cavity implementation hsiehMeanFieldTreatmentVacuum2023, although one could have wished that the change in sampling approach would resolve inaccuracies seen for MF(W) dynamics hoffmannCapturingVacuumFluctuations2019, hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. Regardless, such switching from Wigner sampling to focused sampling allows the number of trajectories, necessary for convergence, to be reduced from (roughly) 500 to 30 for DC-MF dynamics, while a reduction from 5000 to 2500 was found for MF dynamics.
With regard to modeling accuracy, things turn interesting once we apply the aforementioned TDM renormalizations to both methodologies. Fig. 5 compares DC-MF(f1.33) dynamics to DC-MF(W) dynamics where, remarkably, renormalizing the TDM is seen to almost fully remedy the lag known hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023 to occur for the atomic excited-state occupancy within DC-MF(W) dynamics. As such, the applied renormalization yields a further-improved agreement of this method with CISD results, with minor discrepancies occurring exclusively in the reabsorption events. Also included in the comparison is MF(f1.06) dynamics, results of which show qualitative behavior consistent with previous MF(W) dynamics simulations but with a slightly improved relaxation rate hoffmannCapturingVacuumFluctuations2019, hoffmannBenchmarkingSemiclassicalPerturbative2019.
Figure 6: Optical field intensity at $t=500$ a.u. (a), $t=2200$ a.u. (b), and $t=4000$ a.u. (c), which correspond to instances shortly after the first emission, the first reabsorption, and the second reabsorption events, respectively. Note that the optical field is symmetric with respect to reflection through the cavity center, and only the right half of the field intensity is depicted. Shown are results from MF(f1.06) (orange), DC-MF(W) (blue), and DC-MF(f1.33) (purple) dynamics, compared against those from CISD (gray). Each trace was smoothed by a $0.125$ $\mathrm{\mu}$m moving average for ease of demonstration.
@@
As with the atomic occupancies, optical field intensity profiles generated by MF dynamics and DC-MF dynamics are invariant with regard to whether Wigner sampling or focused sampling is applied (see Fig. S3 in the SI), but renormalizing the TDM has a significant effect on DC-MF dynamics. Shown in Fig. 6 are optical field intensity results for MF(f1.06), DC-MF(W), and DC-MF(f1.33) dynamics, compared against CISD results at different time instances. While DC-MF(W) dynamics captures most qualitative features seen in the CISD results, as discussed before hoffmannCapturingVacuumFluctuations2019, hsiehMeanFieldTreatmentVacuum2023, it is again remarkable to see the extent to which the TDM renormalization improves the accuracy of this method to a near-quantitative level. Indeed, DC-MF(f1.33) dynamics reproduces subtle details of the lagging wavefronts that arise upon reabsorption hsiehMeanFieldTreatmentVacuum2023, with differences creeping in only after two reabsorption events. Combined with the favorable results obtained for the half-wavelength cavity, DC-MF(f1.33) dynamics thus presents itself as a method with applicability across cavity lengths. Moreover, the identical values of the applicable TDM rescaling factor $\gamma$ between the short-cavity and long-cavity regimes offer an optimistic prospect regarding the generality of this factor across diverse problems.
@@ 5 Conclusions and Outlook
In conclusion, DC-MF dynamics offers a route to accurate and low-cost simulations of cQED across cavity lengths, provided that focused sampling and TDM renormalizations are invoked. In demonstrating this, we built heavily on recent work hsiehMixedQuantumClassicalDynamics2025 where MF dynamics (without the decoupling scheme) was shown to reproduce full-quantum results for a half-wavelength cavity. Such was achieved under a focused sampling of initial optical field coordinates, in combination with a slight renormalization of the initial optical ZPE. Here, we present a renormalization of the TDMs as an alternative renormalization procedure that not only produces equally accurate results for MF dynamics, but that also allows DC-MF dynamics to produce desirable results. As mentioned, we find that switching from the commonly-used Wigner sampling technique to focused sampling has only a minor effect on Ehrenfest-based mixed quantum–classical modeling in the long-cavity regime, meaning that it fortuitously does not compromise the high accuracy that DC-MF dynamics is known to reach in this regime hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. TDM renormalizations, on the other hand, provide a further source of accuracy improvement for this method, yielding atomic occupancies and optical field intensity profiles in excellent agreement with full-quantum results.
@@
We consistently find focused sampling to dramatically improve the accuracy of both MF and DC-MF dynamics in the short-cavity regime, yet previous applications of this sampling technique were instead primarily concerned with a reduction of computational cost stockFlow1999, bonellaSemiclassicalImplementationMapping2003, bonellaLinearizedPath202005, kimQuantumClassical2008, mandalQuasiDiabatic2018, runesonSpinMapping2019, ananthPathIntegrals2022. While we did not systematically study the computational cost savings afforded by focused sampling in cQED, which we reserve for a future study, we anecdotally found such savings to oftentimes be significant, especially in the short-cavity regime, where a reduction of the necessary number of trajectories was found of three orders of magnitude. In the long-cavity regime, this reduction was much more modest, with a factor of two being typical. Moreover, we find the number of trajectories necessary for convergence to be strongly dependent on the observable under consideration. Notably, transient occupancies of quantum emitters require about 10 to 30 trajectories for DC-MF dynamics under focused sampling, a number much lower than that typically used in mixed quantum–classical simulations of population dynamics. For the optical field intensity, this number raises to values in the range 50 000 to 100 000, which are more typical.
@@
The results presented in this Article help to further establish DC-MF dynamics as a mixed quantum–classical approach for cQED simulations that optimally balances accuracy and computational efficiency. This not only promotes the use of DC-MF dynamics in studying cQED based on model systems, but it also bears relevance to a broader effort to incorporate quantum modeling in established techniques from classical optics mengFrequency2013, sakkoDynamical2014, ruggenthalerFrom2018, jestadtLight2019, sukharevEfficient2023, weightTheory2023, including finite-difference time-domain, for which DC-MF dynamics offers a preferred route. The cost associated with trajectory sampling is very modest, so long as transient matter-based occupancies are being considered, while the evaluation of optical field intensity profiles remains computationally much more challenging.
@@
Against the favorable performance of DC-MF dynamics, there are various aspects of this method that await a deeper rationale. One such aspect is concerned with the form of the decoupling term appearing in Eq. 10. Broadly speaking, this term may be considered representative of a functional of the atomic quantum state, which vanishes when the ground state is occupied while reaching unity for a fully-occupied excited state hsiehMeanFieldTreatmentVacuum2023. Future efforts may unveil the existence of an exact functional, or at least identify improved proxies that further enhance the method’s accuracy. Another aspect that deserves further attention is concerned with the TDM rescaling factors introduced in the present Article. Ultimately, these factors compensate for the tendency of mixed quantum–classical modeling to overshoot the optical occupancy needed to reproduce Rabi oscillations. This in turn is due to the oscillation frequency decaying to zero in the limit $n^{(0)}\rightarrow 0$ hsiehMixedQuantumClassicalDynamics2025. It would be interesting to further investigate this behavior, possibly through an analysis similar to that presented in Ref. hsiehMixedQuantumClassicalDynamics2025. Altogether, such inquiries may provide further guidance for the application of mixed quantum–classical dynamics to the modeling of transient phenomena.
Supporting Information
Methodological details of CISD dynamics and focused sampling results without TDM renormalizations.